A 90-gram wire that is 10 cm long is under tension. When a transverse wave of frequency 890 Hz travels down the wire, it's wavelength is 0.10 and its amplitude is 6.5mm. What is the tension in the wire? What is the mode of this standing wave?
Answer #1

The \; wave \; velocity\; is\; given\; by

wave on string m l

velocity \; of \; wave (v)=f\times \lambda

                                       (v)=890Hz\times .10m

                                          (v)=89 m/s

now \; use\; the \; above\; formula

                     89m/s=\sqrt{\frac{T}{.09/.1}}

                     89m/s=\sqrt{\frac{T}{.9}}

                                   \mathbf{T=7128.9N}\; \; \; (ans)